OpenStudy (fswiney2):

The function for the cost of materials to make a shirt is f(x) = x + 5, where x is the number of shirts. The function for the selling price of those shirts is g(f(x)), where g(x) = 5x + 6. Find the selling price of 18 shirts.

OpenStudy (sachintha):

Plug in the value of \(x\) which is 18 to \(f(x)\) first and the plug the output of it to \(g(x)\). Find \(f(18)\) and use the value of it for \(g(f(18))\). Or, you can simplify the equation and substitute \(x\) = 18 to it. \(g(f(x))=5(x+5)+6=5x+31\)

OpenStudy (radar):

Why not use the selling price function g(x) = 5x = 6 ? g(x) = 5(x) + 6 5(18) + 6 = 90 + 6 = 96 If we are dealing in dollars or pounds, pretty high selling price.